In a pipe that has a cross-sectional area of 0.5 m^2 and a flow rate of 1.5 m^3/s, what is the average speed of the fluid carried by the pipe? *




1
Expert's answer
2021-04-26T18:50:32-0400

Solution:-



So Q = A"\overline v"


Flow rate is the volume of fluid per unit time flowing past a point through the area A. Here the shaded cylinder of fluid flows past point P in a uniform pipe in time t. The volume of the cylinder is Ad and the average velocity is "\bar v = \frac{d}{t}"


So that the flow rate is

"\boxed{Q={Ad\over t }=A\bar v}"


Q(flow rate)= "1.5m^3\/s"


A(cross-sectional area) = 0.5 m2



average speed=


"\boxed{\bar v= \frac{Q}{A}}"


"\boxed{\bar v= \frac{1.5}{0.5}m/s}"


Answer:-

"\boxed{\bar v = 3 m/s}"